# An extension of May’s Theorem to three alternatives: axiomatizing Minimax voting

### v2

@@ 4 Independence of axioms
Case 2: $x$’s Minimax score is 1 greater than the minimal score in $\mathbf{P}$, and $x$ beats the Minimax winner head-to-head. Let $y$ be the Minimax winner and $z$ the alternative that beats $x$ head-to-head in $\mathbf{P}$. By condition (I), the difference between the Minimax scores of $y$ and $z$ in $\mathbf{P}$ is at least 2. Then since $y$’s Minimax score increases by 2 from $\mathbf{P}$ to $\mathbf{P}^{\prime}$, while $z$’s does not decrease, $y$’s Minimax score is no greater than $z$’s Minimax score in $\mathbf{P}^{\prime}$. Hence either $y$ and $z$ have the same Minimax score in $\mathbf{P}^{\prime}$, in which case (I) does not hold for $\mathbf{P}^{\prime}$, or $y$’s Minimax score is smaller than $z$’s, in which case condition (III) does not obtain in $\mathbf{P}^{\prime}$, since $y$ does not beat $x$ head-to-head in $\mathbf{P}^{\prime}$. In either case, $F(\mathbf{P}^{\prime})=Minimax(\mathbf{P}^{\prime})=\{x\}$.∎
@@
and (ii) no voter submits the ranking $yxz$, in which case $F(\mathbf{P})=\{x\}$. Note that we are using ‘$x$’, ‘$y$’, and ‘$z$’ here as variables, so if there is any assignment of the alternatives in $\mathbf{P}$ to the roles $x,y,z$ such that (i) and (ii) holds, then $F$ selects the alternative playing the $x$ role, rather than the Minimax winner. That there are profiles satisfying (i) and (ii), so $F$ differs from Minimax, is shown by the following, where $F(\mathbf{P})=\{x\}$:
@@
Adding a block of all linear orders to the profile above—and hence adding the ranking $yxz$—changes the $F$ winner to be the Minimax winner, $y$, so $F$ violates block preservation. Yet clearly $F$ satisfies anonymity, neutrality, homogeneity, and immunity to spoilers (since $F$ satisfies Condorcet consistency). It remains to prove positive involvement and weak positive responsiveness. Consider an arbitrary profile $\mathbf{P}$.
@@
In this case, we need a lemma: if $a\in F(\mathbf{P})$, $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one voter who ranks $a$ in first place or having some voter move $a$ from last to first place, and as a result condition (i) no longer holds in $\mathbf{P}^{\prime}$ with $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$, then $F(\mathbf{P}^{\prime})=\{a\}$. For suppose $\mathbf{P}$ and $\mathbf{P}^{\prime}$ are as described. Since condition (i) does not hold in $\mathbf{P}^{\prime}$ with $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$, either $0\not\leq Margin_{\mathbf{P}^{\prime}}(c,a)-Margin_{\mathbf{P}^{\prime}}(a,b)$ or $Margin_{\mathbf{P}^{\prime}}(c,a)-Margin_{\mathbf{P}^{\prime}}(a,b)\not<Margin_{\mathbf{P}^{\prime}}(b,c)-Margin_{\mathbf{P}^{\prime}}(c,a)$. But we can rule out the latter possibility, since by the assumption of the case, we have $0\leq Margin_{\mathbf{P}}(c,a)-Margin_{\mathbf{P}}(a,b)<Margin_{\mathbf{P}}(b,c)-Margin_{\mathbf{P}}(c,a)$, and the margin of $a$ over $b$ increases from $\mathbf{P}$ to $\mathbf{P}^{\prime}$ by at least as much as the margin of $b$ over $c$ decreases from $\mathbf{P}$ to $\mathbf{P}^{\prime}$, while the margin of $c$ over $a$ decreases. Thus, we conclude that $0\not\leq Margin_{\mathbf{P}^{\prime}}(c,a)-Margin_{\mathbf{P}^{\prime}}(a,b)$, i.e., $Margin_{\mathbf{P}^{\prime}}(c,a)<Margin_{\mathbf{P}^{\prime}}(a,b)$. Note that since $Margin_{\mathbf{P}}(c,a)<Margin_{\mathbf{P}}(b,c)$ and the margin of $c$ over $a$ decreases by at least as much as the margin of $b$ over $c$ decreases from $\mathbf{P}$ to $\mathbf{P}^{\prime}$, we have $Margin_{\mathbf{P}^{\prime}}(c,a)<Margin_{\mathbf{P}^{\prime}}(b,c)$. Now there are two cases:
@@
Case 2.1.1.1: the margin graph of $\mathbf{P}^{\prime}$ is still as in (i) with $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$. Then since (ii) still holds in $\mathbf{P}^{\prime}$, we still have $F(\mathbf{P})=\{a\}$.
Case 2.1.1.2: the margin graph of $\mathbf{P}^{\prime}$ is no longer as in (i) with $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$. Then by the lemma established above, we again have $F(\mathbf{P})=\{a\}$.
@@
Case 2.1.2.2: $n=m$. Then $F(\mathbf{P})=Minimax(\mathbf{P})=\{a,b\}$. If we add a voter who ranks $b$ in first place or have some voter move $b$ from last to first, then condition (i) still holds for $\mathbf{P}^{\prime}$ with $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$, and condition (ii) still does not hold for $\mathbf{P}^{\prime}$, so $F(\mathbf{P}^{\prime})=Minimax(\mathbf{P}^{\prime})=\{b\}$. Now suppose we add a voter who ranks $a$ in first place or have some voter move $a$ from last to first. Then since $n=m$, in $\mathbf{P}^{\prime}$ the margin of $c$ over $a$ is uniquely smallest among distinct alternatives, so condition (i) no longer holds in $\mathbf{P}^{\prime}$ with $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$. Therefore, by the lemma established above, $F(\mathbf{P}^{\prime})=\{a\}$.
@@
Case 2.2.2: $n<m$, $n<k$, and $k-m\leq m-n$. Since $n<m$ and $n<k$, we have $F(\mathbf{P})=Minimax(\mathbf{P})=\{b\}$. Now suppose $\mathbf{P}^{\prime}$ is obtained by one voter moving $b$ from last to first. Since in $\mathbf{P}^{\prime}$, the margin of $a$ over $b$ is uniquely smallest among distinct alternatives, if $\mathbf{P}^{\prime}$ satisfies (i), it must be under the bijection $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$. Now since in $\mathbf{P}$, we have $k-m\leq m-n$, i.e., $Margin_{\mathbf{P}}(b,c)-Margin_{\mathbf{P}}(c,a)\leq Margin_{\mathbf{P}}(c,a)-Margin_{\mathbf{P}}(a,b)$, we still have $Margin_{\mathbf{P}^{\prime}}(b,c)-Margin_{\mathbf{P}^{\prime}}(c,a)\leq Margin_{\mathbf{P}^{\prime}}(c,a)-Margin_{\mathbf{P}^{\prime}}(a,b)$. Hence (i) does not hold, so $F(\mathbf{P}^{\prime})=Minimax(\mathbf{P}^{\prime})=\{b\}$. On the other hand, suppose $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one voter who ranks $b$ in first place. Again note that since in $\mathbf{P}^{\prime}$, the margin of $a$ over $b$ is uniquely smallest among distinct alternatives, if $\mathbf{P}^{\prime}$ satisfies (i), it must be under the bijection $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$. Now the ranking of the new voter in $\mathbf{P}^{\prime}$ is either $bac$ or $bca$. First, suppose it is $bac$. If (i) does not hold for $\mathbf{P}^{\prime}$, then $F(\mathbf{P}^{\prime})=Minimax(\mathbf{P}^{\prime})=\{b\}$. If it does, then due to the presence of $bac$ in $\mathbf{P}^{\prime}$, (ii) does not hold for $\mathbf{P}^{\prime}$, so $F(\mathbf{P}^{\prime})=Minimax(\mathbf{P}^{\prime})=\{b\}$. Now suppose the new voter’s ranking is $bca$. Then by the assumption of the case, it follows that $Margin_{\mathbf{P}^{\prime}}(b,c)-Margin_{\mathbf{P}^{\prime}}(c,a)\leq Margin_{\mathbf{P}^{\prime}}(c,a)-Margin_{\mathbf{P}^{\prime}}(a,b)$. Hence (i) does not hold, so $F(\mathbf{P}^{\prime})=Minimax(\mathbf{P}^{\prime})=\{b\}$.∎
@@ 5 Beyond three alternatives
The obvious next step is to go beyond three alternatives. One might argue that some of the axioms in Theorem 3.8 ought to be qualified when applied to profiles with four or more alternatives. [^14] See, e.g., the discussion of partial immunity to spoilers in Holliday and Pacuit 2023a, § 4. However, it is certainly of interest to consider the axioms as stated. Table 1 below shows satisfaction/violation of the key axioms in arbitrary profiles by five voting methods (from Schulze 2011, Kemeny 1959, Tideman 1987, and Holliday and Pacuit 2023a, b, respectively) that refine Minimax in three-alternatives profiles. [^15] For example violations, see Pérez 2001, Schulze 2023, and https://github.com/wesholliday/may-minimax. For proofs of satisfaction for Split Cycle, see Holliday and Pacuit 2023a. To see that Kemeny satisfies weak positive responsiveness, we give the argument for Kemeny defined on profiles of linear orders. The argument also works for generalizations of Kemeny to non-linear strict weak orders (Fagin et al. 2006, Zwicker 2018), but we omit the details of these generalizations. Given a profile $\mathbf{P}$, define the Kemeny score of a linear order $L$ of $X(\mathbf{P})$ as the sum over all $(x,y)\in L$ of the number of voters who rank $y$ above $x$ (see Fischer et al. 2016, p. 87). Then $Kemeny(\mathbf{P})$ is the set of alternatives from $X(\mathbf{P})$ that are maximal in some linear order of $X(\mathbf{P})$ with minimal Kemeny score. Now suppose $a\in Kemeny(\mathbf{P})$ and $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by one voter moving $a$ from uniquely last place to uniquely first place. Given $a\in Kemeny(\mathbf{P})$, $a$ is maximal in some linear order $L$ of $X(\mathbf{P})$ with minimal Kemeny score, so for every $b\in X(\mathbf{P})\setminus\{a\}$, we have $(a,b)\in L$. For each such $b$, the number of voters who rank $b$ above $a$ decreases by 1 from $\mathbf{P}$ to $\mathbf{P}^{\prime}$, so each term involving $a$ in the summation for the Kemeny score of $L$ decreases by 1. By contrast, if $L^{\prime}$ does not have $a$ in first place, then the Kemeny score of $L^{\prime}$ cannot decrease as much from $\mathbf{P}$ to $\mathbf{P}^{\prime}$. Hence the orders with minimal Kemeny score in $\mathbf{P}^{\prime}$ all have $a$ in first place, so $Kemeny(\mathbf{P}^{\prime})=\{a\}$, in line with weak positive responsiveness.
@@
Table 1: satisfaction ($\checkmark$) or violation (--) of axioms for arbitrary profiles. All violations are known to happen already for 4 alternatives except for Stable Voting’s violation of weak positive responsiveness, for which the smallest known violation has 8 alternatives. Methods with -- in the third row violate even near immunity to spoilers while methods with $\checkmark$ in the third row satisfy immunity to spoilers.
@@
Do the axioms of Theorem 3.8 characterize the class of refinements of Minimax for more than three alternatives?
Of course, Minimax’s satisfaction of the axioms in Theorem 3.8 in arbitrary profiles must be weighed against its violation—in profiles with four or more alternatives—of other axioms, such as the Condorcet loser criterion (never select as alternative who has negative margins versus every other alternative), satisfied by the other voting methods in Table 1.161616The other methods in Table 1 also satisfy the stronger Smith criterion. The Smith set of a profile $\mathbf{P}$, $Smith(\mathbf{P})$, is the smallest set of alternatives such that every alternative in the set has a positive margin over every alternative outside the set. The Smith criterion is the requirement on voting methods $F$ that $F(\mathbf{P})\subseteq Smith(\mathbf{P})$ for all $\mathbf{P}\in\mathrm{dom}(F)$. We note that the Smith-Minimax method, which chooses the Minimax winners from $\mathbf{P}_{\mid Smith(\mathbf{P})}$, violates positive involvement already for four alternatives (this follows from the impossibility theorem of Holliday 2024).
We already have some hints concerning Question 5.1. Say that a voting method $F$ satisfies ordinal margin invariance if $F(\mathbf{P})=F(\mathbf{P}^{\prime})$ for any profiles $\mathbf{P},\mathbf{P}^{\prime}$ that have the same ordinal margin graph, which is obtained from their margin graphs by replacing the edge weights by the strict weak order of their edges by edge weight. All of the voting methods in Table 1 except for Kemeny satisfy ordinal margin invariance. Now consider the following impossibility theorem from Holliday 2024: already for four alternatives, there is no voting method satisfying positive involvement, the Condorcet winner and loser criteria, ordinal margin invariance, and resolvability. Here resolvability can be replaced by the condition that the voting method selects a unique winner in any profile that is uniquely-weighted, meaning that for all alternatives $a,b,c,d$ such that $a\neq b$, $c\neq d$, $(a,b)\neq(c,d)$, we have $\mathrm{Margin}_{\mathbf{P}}(a,b)\neq\mathrm{Margin}_{\mathbf{P}}(c,d)$ (see Holliday 2024, Lemma 3). Then we can relate the stated impossibility theorem to Question 5.1 as follows.
@@
Any voting method satisfying the axioms of Theorem 3.8 and ordinal margin invariance for four alternatives must violate the Condorcet loser criterion for four alternatives.
@@
Thus, $F$ selects a unique winner in any uniquely-weighted profile. Moreover, we showed in the proof of Theorem 3.8 that if $F$ satisfies the axioms of Theorem 3.8, then $F$ satisfies the Condorcet winner criterion. Thus, by the main result of Holliday 2024, $F$ violates the Condorcet loser criterion for four alternatives.∎

### v4

@@ 4 Independence of axioms
Case 2: $x$’s Minimax score is 1 greater than the minimal score in $\mathbf{P}$, and $x$ beats the Minimax winner head-to-head. Let $y$ be the Minimax winner and $z$ the alternative that beats $x$ head-to-head in $\mathbf{P}$. By condition (I), the difference between the Minimax scores of $y$ and $z$ in $\mathbf{P}$ is at least 2. Then since $y$’s Minimax score increases by only 2 from $\mathbf{P}$ to $\mathbf{P}^{\prime}$, while $z$’s does not decrease, $y$’s Minimax score is no greater than $z$’s Minimax score in $\mathbf{P}^{\prime}$. Hence either $y$ and $z$ have the same Minimax score in $\mathbf{P}^{\prime}$, in which case condition (I) does not hold for $\mathbf{P}^{\prime}$, or $y$’s Minimax score is smaller than $z$’s, in which case condition (III) does not obtain in $\mathbf{P}^{\prime}$, since $y$ does not beat $x$ head-to-head in $\mathbf{P}^{\prime}$. In either case, $F(\mathbf{P}^{\prime})=Minimax(\mathbf{P}^{\prime})=\{x\}$.∎
@@
and (ii) no voter submits the ranking $yxz$, in which case $F(\mathbf{P})=\{x\}$. Note that we are using ‘$x$’, ‘$y$’, and ‘$z$’ here as variables, so if there is any assignment of the alternatives in $\mathbf{P}$ to the roles $x,y,z$ such that (i) and (ii) hold, then $F$ selects the alternative playing the $x$ role, rather than the Minimax winner. That there are profiles satisfying (i) and (ii), so $F$ differs from Minimax, is shown by the following, where $F(\mathbf{P})=\{x\}$:
@@
Adding a block of all linear orders to the profile above—and hence adding the ranking $yxz$—changes the $F$ winner to be the Minimax winner, $y$, so $F$ violates block preservation. Yet clearly $F$ satisfies anonymity, neutrality, homogeneity, and immunity to spoilers (since $F$ agrees with Minimax in all of the profiles satisfying the assumptions in the definition of immunity to spoilers). It remains to prove positive involvement and weak positive responsiveness. Consider an arbitrary profile $\mathbf{P}$.
@@
In this case, we need a lemma: if $a\in F(\mathbf{P})$, $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one voter who ranks $a$ in first place or having some voter move $a$ from last to first place, and as a result condition (i) no longer holds in $\mathbf{P}^{\prime}$ with $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$, then $F(\mathbf{P}^{\prime})=\{a\}$. To see this, suppose $\mathbf{P}$ and $\mathbf{P}^{\prime}$ are as described. Since condition (i) does not hold in $\mathbf{P}^{\prime}$ with $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$, either $0\not\leq Margin_{\mathbf{P}^{\prime}}(c,a)-Margin_{\mathbf{P}^{\prime}}(a,b)$ or $Margin_{\mathbf{P}^{\prime}}(c,a)-Margin_{\mathbf{P}^{\prime}}(a,b)\not<Margin_{\mathbf{P}^{\prime}}(b,c)-Margin_{\mathbf{P}^{\prime}}(c,a)$. But we can rule out the latter possibility, since by the assumption of the case, we have $0\leq Margin_{\mathbf{P}}(c,a)-Margin_{\mathbf{P}}(a,b)<Margin_{\mathbf{P}}(b,c)-Margin_{\mathbf{P}}(c,a)$, and the margin of $a$ over $b$ increases from $\mathbf{P}$ to $\mathbf{P}^{\prime}$ by at least as much as the margin of $b$ over $c$ decreases from $\mathbf{P}$ to $\mathbf{P}^{\prime}$, while the margin of $c$ over $a$ decreases. Thus, we conclude that $0\not\leq Margin_{\mathbf{P}^{\prime}}(c,a)-Margin_{\mathbf{P}^{\prime}}(a,b)$, i.e., $Margin_{\mathbf{P}^{\prime}}(c,a)<Margin_{\mathbf{P}^{\prime}}(a,b)$. Note that since $Margin_{\mathbf{P}}(c,a)<Margin_{\mathbf{P}}(b,c)$ and the margin of $c$ over $a$ decreases by at least as much as the margin of $b$ over $c$ decreases from $\mathbf{P}$ to $\mathbf{P}^{\prime}$, we have $Margin_{\mathbf{P}^{\prime}}(c,a)<Margin_{\mathbf{P}^{\prime}}(b,c)$. Now there are two cases:
@@
Case 2.1.1.1: the margin graph of $\mathbf{P}^{\prime}$ is still as in (i) with $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$. Then since (ii) still holds in $\mathbf{P}^{\prime}$, we still have $F(\mathbf{P}^{\prime})=\{a\}$.
Case 2.1.1.2: the margin graph of $\mathbf{P}^{\prime}$ is no longer as in (i) with $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$. Then by the lemma established above, we again have $F(\mathbf{P}^{\prime})=\{a\}$.
@@
Case 2.1.2.2: $n=m$. Then $F(\mathbf{P})=Minimax(\mathbf{P})=\{a,b\}$. If we add a voter who ranks $b$ in first place or have some voter move $b$ from last to first, then condition (i) still holds for $\mathbf{P}^{\prime}$ with $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$, and condition (ii) still does not hold for $\mathbf{P}^{\prime}$, so $F(\mathbf{P}^{\prime})=Minimax(\mathbf{P}^{\prime})=\{b\}$. Now suppose we add a voter who ranks $a$ in first place or have some voter move $a$ from last to first. Then since $n=m$, in $\mathbf{P}^{\prime}$ the margin of $c$ over $a$ is smaller than that of $a$ over $b$ and that of $b$ over $c$, so condition (i) no longer holds in $\mathbf{P}^{\prime}$ with $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$. Therefore, by the lemma established above, $F(\mathbf{P}^{\prime})=\{a\}$.
@@
Case 2.2.2: $n<m$, $n<k$, and $k-m\leq m-n$. Since $n<m$ and $n<k$, we have $F(\mathbf{P})=Minimax(\mathbf{P})=\{b\}$. Now suppose $\mathbf{P}^{\prime}$ is obtained by one voter moving $b$ from last to first. Since in $\mathbf{P}^{\prime}$, the margin of $a$ over $b$ is smaller than that of $b$ over $c$ and that of $c$ over $a$, if $\mathbf{P}^{\prime}$ satisfies (i), it must be under the bijection $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$. Now since in $\mathbf{P}$, we have $k-m\leq m-n$, i.e., $Margin_{\mathbf{P}}(b,c)-Margin_{\mathbf{P}}(c,a)\leq Margin_{\mathbf{P}}(c,a)-Margin_{\mathbf{P}}(a,b)$, we still have $Margin_{\mathbf{P}^{\prime}}(b,c)-Margin_{\mathbf{P}^{\prime}}(c,a)\leq Margin_{\mathbf{P}^{\prime}}(c,a)-Margin_{\mathbf{P}^{\prime}}(a,b)$. Hence (i) does not hold, so $F(\mathbf{P}^{\prime})=Minimax(\mathbf{P}^{\prime})=\{b\}$. On the other hand, suppose $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one voter who ranks $b$ in first place. Again note that since in $\mathbf{P}^{\prime}$, the margin of $a$ over $b$ is smaller than that of $b$ over $c$ and that of $c$ over $a$, if $\mathbf{P}^{\prime}$ satisfies (i), it must be under the bijection $a\mapsto x$, $b\mapsto y$, and $c\mapsto z$. Now the ranking of the new voter in $\mathbf{P}^{\prime}$ is either $bac$ or $bca$. First, suppose it is $bac$. If (i) does not hold for $\mathbf{P}^{\prime}$, then $F(\mathbf{P}^{\prime})=Minimax(\mathbf{P}^{\prime})=\{b\}$. If it does, then due to the presence of $bac$ in $\mathbf{P}^{\prime}$, (ii) does not hold for $\mathbf{P}^{\prime}$, so $F(\mathbf{P}^{\prime})=Minimax(\mathbf{P}^{\prime})=\{b\}$. Now suppose the new voter’s ranking is $bca$. Then by the assumption of the case, it follows that $Margin_{\mathbf{P}^{\prime}}(b,c)-Margin_{\mathbf{P}^{\prime}}(c,a)\leq Margin_{\mathbf{P}^{\prime}}(c,a)-Margin_{\mathbf{P}^{\prime}}(a,b)$. Hence (i) does not hold, so $F(\mathbf{P}^{\prime})=Minimax(\mathbf{P}^{\prime})=\{b\}$.∎
@@ 5 Beyond three alternatives
The obvious next step is to go beyond three alternatives. One might argue that some of the axioms in Theorem 3.8 ought to be qualified when applied to profiles with four or more alternatives. [^15] See, e.g., the discussion of partial immunity to spoilers in Holliday and Pacuit 2023a, § 4. However, it is certainly of interest to consider the axioms as stated. Table 1 below shows satisfaction/violation of the key axioms in arbitrary profiles by five voting methods (from Schulze 2011, Kemeny 1959, Tideman 1987, and Holliday and Pacuit 2023a, b, respectively) that refine Minimax in three-alternative profiles. [^16] For examples of violations, see Pérez 2001, Schulze 2023, and https://github.com/wesholliday/may-minimax. For proofs of satisfaction for Split Cycle, see Holliday and Pacuit 2023a. To see that Kemeny satisfies weak positive responsiveness, we give the argument for Kemeny defined on profiles of linear orders. The argument also works for generalizations of Kemeny to non-linear strict weak orders (Fagin et al. 2006, Zwicker 2018), but we omit the details of these generalizations. Given a profile $\mathbf{P}$, define the Kemeny score of a linear order $L$ of $X(\mathbf{P})$ as the sum over all $(x,y)\in L$ of the number of voters who rank $y$ above $x$ (see Fischer et al. 2016, p. 87). Then $Kemeny(\mathbf{P})$ is the set of alternatives from $X(\mathbf{P})$ that are maximal in some linear order on $X(\mathbf{P})$ with minimal Kemeny score. Now suppose $a\in Kemeny(\mathbf{P})$ and $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by one voter moving $a$ from uniquely last place to uniquely first place. Given $a\in Kemeny(\mathbf{P})$, $a$ is maximal in some linear order $L$ of $X(\mathbf{P})$ with minimal Kemeny score, so for every $b\in X(\mathbf{P})\setminus\{a\}$, we have $(a,b)\in L$. For each such $b$, the number of voters who rank $b$ above $a$ decreases by 1 from $\mathbf{P}$ to $\mathbf{P}^{\prime}$, so each term involving $a$ in the summation for the Kemeny score of $L$ decreases by 1. By contrast, if $L^{\prime}$ does not have $a$ in first place, then the Kemeny score of $L^{\prime}$ cannot decrease as much from $\mathbf{P}$ to $\mathbf{P}^{\prime}$. Hence the orders with minimal Kemeny score in $\mathbf{P}^{\prime}$ all have $a$ in first place, so $Kemeny(\mathbf{P}^{\prime})=\{a\}$, in line with weak positive responsiveness. Note, by way of contrast, that all proper scoring rules (see Zwicker 2016, Def. 2.9) satisfy weak positive responsiveness and positive involvement but violate (near) immunity to spoilers already for three alternatives.
@@
Table 1: satisfaction ($\checkmark$) or violation (--) of axioms for arbitrary profiles. All violations are known to happen already for 4 alternatives except for Beat Path’s violation of immunity to spoilers, which requires 5 alternatives, and Stable Voting’s violation of weak positive responsiveness, for which the smallest known violation has 8 alternatives. Methods with -- in the third row violate even near immunity to spoilers while methods with $\checkmark$ in the third row satisfy immunity to spoilers.
@@
Do the axioms of Theorem 3.8 characterize a class of refinements of Minimax for more than three alternatives?
Of course, Minimax’s satisfaction of the axioms in Theorem 3.8 in arbitrary profiles must be weighed against its violation—in profiles with four or more alternatives—of other axioms, such as the Condorcet loser criterion (never select an alternative that has negative margins versus every other alternative), satisfied by the other voting methods in Table 1.171717The other methods in Table 1 also satisfy the stronger Smith criterion. The Smith set of a profile $\mathbf{P}$, $Smith(\mathbf{P})$, is the smallest set of alternatives such that every alternative in the set has a positive margin over every alternative outside the set. The Smith criterion is the requirement on voting methods $F$ that $F(\mathbf{P})\subseteq Smith(\mathbf{P})$ for all $\mathbf{P}\in\mathrm{dom}(F)$. We note that the Smith-Minimax method, which chooses the Minimax winners from $\mathbf{P}_{\mid Smith(\mathbf{P})}$, violates positive involvement already for four alternatives (this follows from the impossibility theorem of Holliday 2024).
We already have some hints concerning Question 5.1. Say that a voting method $F$ satisfies ordinal margin invariance if $F(\mathbf{P})=F(\mathbf{P}^{\prime})$ for any profiles $\mathbf{P},\mathbf{P}^{\prime}$ that have the same ordinal margin graph, which is obtained from their margin graphs by replacing the edge weights by the strict weak order of their edges by edge weight. Ordinal margin invariance implies several of our previous axioms, namely anonymity, homogeneity, and block preservation. Moreover, all of the voting methods in Table 1 except for Kemeny satisfy ordinal margin invariance. Now consider the following impossibility theorem from Holliday 2024: already for four alternatives, there is no voting method satisfying positive involvement, the Condorcet winner and loser criteria, ordinal margin invariance, and resolvability. Here resolvability can be replaced by the condition that the voting method selects a unique winner in any profile that is uniquely weighted, meaning that for all alternatives $a,b,c,d$ such that $a\neq b$, $c\neq d$, $(a,b)\neq(c,d)$, we have $\mathrm{Margin}_{\mathbf{P}}(a,b)\neq\mathrm{Margin}_{\mathbf{P}}(c,d)$ (see Holliday 2024, Lemma 3). Then we can relate the stated impossibility theorem to Question 5.1 as follows.
@@
Let $F$ be a voting method on the domain of profiles with up to four alternatives satisfying neutrality, weak positive responsiveness, positive involvement, near immunity to spoilers, and ordinal margin invariance. Then $F$ violates the Condorcet loser criterion.
@@
Thus, $F$ selects a unique winner in any uniquely-weighted profile. Moreover, the axioms of the current proposition imply the axioms in Theorem 3.8, so by that theorem, $F$ satisfies the Condorcet winner criterion for profiles up to three alternatives. To prove that $F$ satisfies the Condorcet winner criterion for profiles of four alternatives, we reason exactly as in the proof of the first paragraph of Theorem 3.8 except that we invoke Theorem 3.8 itself to prove $F(\mathbf{P}^{\prime\prime}_{-b})=\{c\}$. Thus, $F$ satisfies the Condorcet winner criterion. But then by the main result of Holliday 2024, $F$ violates the Condorcet loser criterion.∎